### v1

# Inferring parameters and reconstruction of two-dimensional turbulent flows with physics-informed neural networks

Abstract
Solving inverse problems, which means obtaining model parameters from observed data, using conventional computational fluid dynamics solvers is prohibitively expensive. Here we employ machine learning algorithms to overcome the challenge. As an example, we consider a moderately turbulent fluid flow, excited by a stationary force and described by a two-dimensional Navier-Stokes equation with linear bottom friction. Given sparse and probably noisy data for the velocity and the general form of the model, we reconstruct the dense velocity and pressure fields in the observation domain, infer the driving force, and determine the unknown fluid viscosity and friction coefficient. Our approach involves training a physics-informed neural network by minimizing the loss function, which penalizes deviations from the provided data and violations of the Navier-Stokes equation. The suggested technique extracts additional information from experimental and numerical observations, potentially enhancing the capabilities of particle image/tracking velocimetry.
Extracting information and reconstructing a flow field from sparse or noisy measurements is important yet difficult in many applications. Solving the problem requires additional knowledge to compensate for imperfections in the data. A well-constructed physical model of the phenomenon can serve in this capacity. However, traditional computational fluid dynamics (CFD) models are ill-suited for these purposes due to the computational complexity of data assimilation methods [1], especially for turbulent flows. Recent advances in machine learning (ML) open up new opportunities for addressing these challenges [2, 3, 4, 5].
In this work, we will focus on physics-informed neural networks (PINNs) [6], which provide a framework for merging data with physical laws expressed as partial differential equations (PDEs). Previously, PINNs were applied to uncover near-wall blood flow from sparse data [7], for super-resolution and denoising applications [8, 9, 10], to reconstruct velocity and pressure fields from flow visualizations [11, 12] and particle image/tracking velocimetry (PIV/PTV) data [13, 14], to simulate turbulence [15, 16, 17], just to name a few. A more comprehensive bibliography can be found in recent reviews [3, 4, 18, 5]. Here we implement PINN to solve inverse problem inspired by experimental studies of forced turbulent flows in thin conducting fluid layers [19, 20, 21, 22, 23, 24]. Our goal is to show that when the physical model is taken into account, the suggested technique has the ability to extract more information from conventional flow observations.
Let us start by defining the problem. We consider two-dimensional incompressible fluid flow described by the forced Navier-Stokes equation with linear bottom friction
$\partial_{t}\bm{v}+(\bm{v}\cdot\nabla)\bm{v}=-\nabla p-\alpha\bm{v}+\nu\nabla^{2}\bm{v}+\bm{f},$ (1)
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\alpha$ is the friction coefficient, and $\nu$ is the kinematic viscosity. The external force $\bm{f}$ is assumed to be stationary in time and divergence-free in space. We are interested in a moderately turbulent regime, when the nonlinear term in the Navier-Stokes equation plays a substantial role. Then, despite the stationary nature of the external forcing, the fluid flow is characterized by strong spatiotemporal variations. Now suppose that someone measures the velocity field in a certain region (possibly with open boundaries) and makes observations for several times longer than the typical time of flow fluctuations. In what follows, we will demonstrate how, using ML techniques, one can not only reconstruct the dense velocity $\bm{v}(\bm{r},t)$ and pressure $p(\bm{r},t)$ fields (super-resolution application), but also establish the driving force $\bm{f}(\bm{r})$ and unknown fluid viscosity $\nu$ and bottom friction $\alpha$, based only on potentially sparse and noisy velocity field data. Let us emphasize that the developed approach does not require knowledge of the boundary conditions at the bounds of the observation area.
Data generation. Data for this study are obtained by integrating (1) using the GeophysicalFlows.jl pseudospectral code [25], which has recently been successfully applied to model two-dimensional turbulence [26, 27, 28]. The code can be executed on the GPU, resulting in high computational performance. The aliasing errors are removed with the two-third rule. The simulation domain is a doubly periodic box of size $2\pi\times 2\pi$. The boundary conditions and form of the domain are only relevant to the data production process; the rest of the study is generic in this regard.
The observation area is smaller and has the size of $L\times L$ with $L=\pi$, see Fig. 1a. Thus, fluid flows in and out of the observation zone and the movement within it depends on the unseen surroundings. The system parameters $\nu=0.01$ and $\alpha=0.1$ are chosen to be consistent with the experimental studies. The external forcing has the Kolmogorov form
$f_{x}=f_{0}\sin(k_{f}y),\quad f_{y}=0,$ (2)
with $f_{0}=10$ and $k_{f}=5$. We start from an arbitrary random initial condition, and after a transient process, the flow reaches a statistical steady-state. Once stationary, we saved data at $\Delta t=0.02$ intervals for $T=4$ units, covering several turn-over times of flow fluctuations. The spatial resolution is $512^{2}$ for the entire computational domain and $256^{2}$ for the observation zone.
Fig. 1a shows snapshot of the vorticity field $\omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ in the statistical steady-state. The flow is chaotic, with no obvious imprint of an external forcing. Since there are two dissipation mechanisms – viscosity and bottom friction – the flow state can be characterized by two dimensionless parameters [29]: the Reynolds number $Re=UL/\nu\approx 1.3\times 10^{3}$ and $Rh=U/(\alpha L)\approx 13$, which were computed using the root mean square velocity $U$ and the size $L$ of the observation domain. The energy spectrum is presented in Fig. 1b. It indicates that fluid flow is mainly determined by modes with wave vectors $k\lesssim k_{f}$.
Figure 1: Vorticity field (a) and energy spectrum (b) in a statistical steady-state. The shaded area shows the observation zone.
Next, we assume that measurements reveal the velocity field $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ at $N_{data}$ points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, which are randomly scattered within the observation area and time interval of length $T$. These data imitate PIV/PTV measurements and will be used for subsequent neural network training. To study the influence of measurement noise, we will distort the velocity as $v_{x,y}^{d}\rightarrow(1+\eta)v_{x,y}^{d}$, where $\eta\sim\mathcal{N}(0,\varepsilon^{2})$ is a normally distributed random variable and the standard deviation $\varepsilon$ controls the noise level. We set $N_{data}=3\times 10^{4}$, which is around $0.2\%$ of the total data and thus represents sparse measurements.
Physics-Informed Neural Network. In our study, we utilize two independent neural networks (NNs) that are trained simultaneously, see Fig. 2a. The first NN approximates the velocity and pressure at a particular position and time. It has three input neurons and three output neurons, $(v_{x},v_{y},p)=\mathcal{F}_{\vec{\theta}_{1}}(x,y,t)$. The second NN is intended to predict stationary force. As a result, it contains only two input neurons and two output neurons, $(f_{x},f_{y})=\mathcal{G}_{\vec{\theta}_{2}}(x,y)$. The parameters $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$ denote the trainable variables for NNs.
Figure 2: (a) Schematic of the neural networks and (b) adopted residual network architecture.
Figure 3: Comparison between DNS data (ground truth) and PINN predictions for model $A_{1}$ (with clean data).
NNs must be sufficiently complicated to reproduce patterns that are essential for the considered turbulent flow. Universal approximation theorem ensures that this is possible [30, 31]. In the original paper [6], the authors used a fully connected NN architecture, but here we implement a residual design [32] for both NNs as we found that it performs better in agreement with the results reported earlier [33]. The adopted architecture is shown in Fig. 2b. The parameters $N_{l}$ and $N_{n}$ represent the number of hidden layers and neurons in each hidden layer, respectively. The output $\bm{Y}_{k}$ of $k$-layer is given by
$\bm{Y}_{2j+1}=\sigma\left(\hat{W}_{2j+1}\bm{Y}_{2j}+\bm{b}_{2j+1}+\bm{Y}_{2j-1}\right),\quad j\geq 0,$
$\bm{Y}_{2j}=\sigma\left(\hat{W}_{2j}\bm{Y}_{2j-1}+\bm{b}_{2j}\right),\quad j\geq 1,$ (3)
$\bm{Y}_{N_{l}+1}=\hat{W}_{N_{l}+1}\bm{Y}_{N_{l}}+\bm{b}_{N_{l}+1},$
where $\bm{Y}_{-1}\equiv\bm{0}$ for convenience of notation, $\bm{Y}_{0}$ corresponds to the input layer, $\bm{Y}_{N_{l}+1}$ denotes the output layer, $\hat{W}$ and $\bm{b}$ are weights and biases (trainable parameters that comprise $\vec{\theta}$), and $\sigma$ is the nonlinear activation function applied element-wise. We use a tanh activation function because it gives a reasonable trade-off between training time and resulting accuracy [13].
In general, increasing the size of NNs can reduce the prediction error, but only when the NNs are given enough data [13]. Here we set $N_{l}=7$ and $N_{n}=250$ for $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$, and $N_{l}=5$ and $N_{n}=30$ for $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$. This choice of parameters is consistent with $N_{data}=3\times 10^{4}$. Note that $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$ is much simpler than $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$ because it is not time dependent and the external forcing usually does not have a complex spatial dependency.
Training NNs involves minimizing a loss function consisting of two terms. The first term represents the deviation between predicted and measured velocity
$\mathcal{L}_{data}=\dfrac{1}{N_{data}U^{2}}\sum_{n=1}^{N_{data}}|\bm{v}(\bm{r}_{n}^{d},t_{n}^{d})-\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})|^{2},$ (4)
where $U^{2}=\sum_{n}\bm{v}_{d}^{2}(\bm{r}_{n}^{d},t_{n}^{d})/N_{data}$ is measurement-based estimate of mean squared velocity. The second term penalizes deviations of predictions from the Navier-Stokes equation and incompressibility conditions
$\mathcal{L}_{eq}=\dfrac{1}{N_{eq}}\sum_{n=1}^{N_{eq}}\left[\bm{e}_{1}^{2}(\bm{r}_{n}^{e},t_{n}^{e})+e_{2}^{2}(\bm{r}_{n}^{e},t_{n}^{e})+e_{3}^{2}(\bm{r}_{n}^{e})\right],$ (5)
where $\{\bm{r}_{n}^{e},t_{n}^{e}\}_{n=1}^{N_{eq}}$ are the collocation points at which equation loss is calculated and
$\bm{e}_{1}=\left(\partial_{t}\bm{v}+(\bm{v}\nabla)\bm{v}+\nabla p+\alpha\bm{v}-\nu\nabla^{2}\bm{v}-\bm{f}\right)L/U^{2},\;$ (6)
$e_{2}=\left(\partial_{x}v_{x}+\partial_{y}v_{y}\right)L/U,$ (7)
$e_{3}=\left(\partial_{x}f_{x}+\partial_{y}f_{y}\right)L^{2}/U^{2}.$ (8)
Computing the right-hand sides involves differentiating the outputs of NNs with respect to input neurons using automatic differentiation [34]. The number and location of collocation points are not limited by measurements and can be chosen arbitrarily. We set $N_{eq}=3N_{data}$, with one-third of these points corresponding to data points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, and the rest are generated randomly. Note that instead of requiring the incompressibility conditions to be satisfied, the NNs can be modified to predict the relevant stream functions [6, 16]. However, in this case, calculating the terms in the Navier-Stokes equation involves computation of higher-order derivatives, which significantly slows down the training process.
Finally, the total loss function is given by
$\mathcal{L}=\mathcal{L}_{data}+\beta\mathcal{L}_{eq},$ (9)
where $\beta$ is a weighting coefficient used to balance two terms. Its value can be adaptively changed during training [15], but here we use a simpler strategy in which the weight is fixed. Empirically, we have found that $\beta=0.02$ is well suited for this type of problem.
Minimization of the loss function (9) is performed by a gradient-based optimizer that adjusts network and model parameters. As the bottom friction and fluid viscosity are positive, it is convenient to represent them in the form $\alpha=e^{\lambda_{1}}$ and $\nu=e^{\lambda_{2}}$. So, the training variables are $\{\vec{\theta}_{1},\vec{\theta}_{2},\lambda_{1},\lambda_{2}\}$. First, we train the model for $10^{5}$ epochs using Adam optimizer [35] with learning rate of $\mu=10^{-3}$. To reduce computation time, we have implemented a mixed precision technique [36]. Then we use L-BFGS-B optimizer [37] until the tolerance reaches machine precision. The training is carried out on an NVIDIA A100 card. The code is written using the PyTorch library [38] and is openly available at [39].
[…]
$\Delta_{Q}=\left\langle\dfrac{||Q-Q_{DNS}||_{2}}{||Q_{DNS}||_{2}}\right\rangle_{T},$ (10)
where $Q$ denotes the physical quantity of interest predicted by the PINN, and $Q_{DNS}$ is its exact value. Table 1 shows the results for velocity, pressure and external force, as well as the relative errors in determining the bottom friction $\alpha$ and fluid viscosity $\nu$. The developed method consistently reconstructs the velocity field in the observation region with an accuracy of about $0.2\%$, while the model parameters $\alpha$ and $\nu$ are determined with an error of several percent. Relative errors in predicting the pressure field and external force are larger, with greater dispersion and an obvious correlation between them. The average for both is about $8\%$, and the standard deviation is close to $5\%$. This behavior is due to the lack of measurements for these quantities, so the model extracts them solely from the analysis of physical laws.
Table 1: Relative $L_{2}$ errors in predicting velocity, pressure, external force, bottom friction and fluid viscosity (with clean data).
model $\Delta_{\bm{v}},\%$ $\Delta_{p-\bar{p}},\%$ $\Delta_{\bm{f}},\%$ $\Delta_{\alpha},\%$ $\Delta_{\nu},\%$
$A_{0}$ $0.22$ $5.52$ $6.09$ $3.66$ $1.42$
$A_{1}$ $0.21$ $4.21$ $4.02$ $4.27$ $1.56$
$A_{2}$ $0.20$ $3.11$ $4.10$ $2.83$ $1.41$
$A_{3}$ $0.23$ $17.34$ $16.58$ $4.36$ $1.50$
$A_{4}$ $0.20$ $2.13$ $2.28$ $3.22$ $1.26$
$A_{5}$ $0.24$ $8.34$ $8.39$ $3.84$ $1.76$
$A_{6}$ $0.22$ $10.46$ $10.17$ $4.02$ $1.47$
$A_{7}$ $0.21$ $17.37$ $16.59$ $4.21$ $1.53$
$A_{8}$ $0.21$ $3.85$ $3.66$ $4.07$ $1.65$
$A_{9}$ $0.23$ $7.48$ $7.25$ $4.10$ $1.59$
Next, we repeat the training, but now the models receive noisy measurement data $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$. Table 2 reports the dependence of the average values and standard deviations for prediction errors on the noise level, obtained by averaging over ten models in each case. The developed algorithm is robust to small noise ($\leq 1\%$) in the input data. Moreover, the error for the reconstructed velocity field in this case is less than the noise level. Thus, PINN is able to correct for small noise based on the knowledge that the velocity field must satisfy the Navier-Stokes equation and incompressibility condition. As the noise level increases ($\geq 2\%$), the performance of the trained models gradually decreases. Since for a turbulent flow the dissipative terms make a small contribution to the Navier-Stokes equation, the model quickly loses the ability to predict the bottom friction $\alpha$ and fluid viscosity $\nu$. At the same time, the accuracy of the velocity field determination remains comparable to the noise level. Errors in predicting the pressure field and external force are larger and amount to tens of percent.
Table 2: Dependence of mean values and standard deviations for prediction errors on the noise level.
noise $\overline{\Delta}_{\bm{v}},\%$ $\overline{\Delta}_{p-\bar{p}},\%$ $\overline{\Delta}_{\bm{f}},\%$ $\overline{\Delta}_{\alpha},\%$ $\overline{\Delta}_{\nu},\%$
$0\%$ $0.22\pm 0.01$ $8.0\pm 5.6$ $7.9\pm 5.2$ $3.9\pm 0.5$ $1.5\pm 0.1$
$0.5\%$ $0.28\pm 0.01$ $8.9\pm 5.4$ $8.6\pm 5.1$ $3.1\pm 0.9$ $1.2\pm 0.2$
$1\%$ $0.50\pm 0.03$ $7.8\pm 5.9$ $7.6\pm 5.6$ $2.2\pm 1.8$ $0.7\pm 0.5$
$2\%$ $2.39\pm 0.06$ $12.9\pm 5.1$ $12.2\pm 5.3$ $41.4\pm 4.4$ $12.3\pm 0.9$
$5\%$ $5.84\pm 0.28$ $19.1\pm 5.5$ $17.9\pm 6.0$ $188\pm 26$ $56.1\pm 2.0$
In conclusion, we demonstrated that PINN can integrate measurement data with an accurate physical model to extract additional information from ordinary flow observations. The developed approach was based on joint training of two independent NNs, each of which took into account the properties of individual terms of the Navier-Stokes equation, encoding them in its own architecture. The results obtained for turbulent flow showed good accuracy for super-resolution and inference applications and demonstrated moderate robustness to measurement noise. The suggested technique may be useful for analyzing observations (e.g., of the ocean or atmosphere), when it is difficult to carry out additional studies aimed at establishing the parameters (e.g., turbulent viscosity) of the physical model.
Acknowledgements.
The work of VP and IN was supported by the Ministry of Science and Higher Education of the Russian Federation within the State Assignment No. FFWR-2024-0017 of Landau Institute for Theoretical Physics. VP also acknowledges support from the Foundation for the Advancement of Theoretical Physics and Mathematics ”BASIS”, Project No. 22-1-3-24-1. The authors are grateful to the Landau Institute for Theoretical Physics for providing computing resources.