### v2

# Enhancing capabilities of particle image/tracking velocimetry with physics-informed neural networks

Abstract
Solving inverse problems, which means obtaining model parameters from observed data, using conventional computational fluid dynamics solvers is prohibitively expensive. Here we employ machine learning algorithms to overcome the challenge. As an example, we consider a moderately turbulent fluid flow, excited by a stationary force and described by a two-dimensional Navier-Stokes equation with linear bottom friction. Given sparse and probably noisy data for the velocity and the general form of the model, we reconstruct the dense velocity and pressure fields in the observation domain, infer the driving force, and determine the unknown fluid viscosity and friction coefficient. Our approach involves training a physics-informed neural network by minimizing the loss function, which penalizes deviations from the provided data and violations of the Navier-Stokes equation. The suggested technique extracts additional information from experimental and numerical observations, potentially enhancing the capabilities of particle image/tracking velocimetry.
I Introduction
Extracting information and reconstructing a flow field from sparse or noisy measurements is important yet difficult in many applications. Solving the problem requires additional knowledge to compensate for imperfections in the data. A well-constructed physical model of the phenomenon can serve in this capacity. However, traditional computational fluid dynamics (CFD) models are ill-suited for these purposes due to the computational complexity of data assimilation methods Asch, Bocquet, and Nodet (2016), especially for turbulent flows. Recent advances in machine learning (ML) open up new opportunities for addressing these challenges Brunton, Noack, and Koumoutsakos (2020); Cai et al. (2021a); Karniadakis et al. (2021); Sharma et al. (2023).
In this work, we focus on physics-informed neural networks (PINNs) Lagaris, Likas, and Fotiadis (1998); Raissi, Perdikaris, and Karniadakis (2019), which provide a framework for merging data with physical laws expressed as partial differential equations (PDEs). Previously, PINNs were applied to uncover near-wall blood flow from sparse data Arzani, Wang, and D’Souza (2021), for super-resolution and denoising applications Fathi et al. (2020); Gao, Sun, and Wang (2021); Eivazi and Vinuesa (2022), to reconstruct velocity and pressure fields from flow visualizations Raissi, Yazdani, and Karniadakis (2020); Cai et al. (2021b); Clark Di Leoni et al. (2023a) and particle image/tracking velocimetry (PIV/PTV) data Wang, Liu, and Wang (2022); Clark Di Leoni et al. (2023b), to simulate turbulence Jin et al. (2021); Kag, Seshasayanan, and Gopinath (2022); Eivazi et al. (2022), just to name a few. A more comprehensive bibliography can be found in recent reviews Cai et al. (2021a); Karniadakis et al. (2021); Cuomo et al. (2022); Sharma et al. (2023). Here we implement PINN to solve inverse problem inspired by experimental studies of forced turbulent flows in thin conducting fluid layers Sommeria (1986); Paret and Tabeling (1998); Boffetta et al. (2005); Xia, Shats, and Falkovich (2009); Orlov, Brazhnikov, and Levchenko (2018); Fang and Ouellette (2021). Our goal is to show that when the physical model is taken into account, the suggested technique has the ability to extract more information from conventional flow observations.
We consider two-dimensional incompressible fluid flow described by the forced Navier-Stokes equation with linear bottom friction
$\partial_{t}\bm{v}+(\bm{v}\cdot\nabla)\bm{v}=-\nabla p-\alpha\bm{v}+\nu\nabla^{2}\bm{v}+\bm{f},$ (1)
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\alpha$ is the friction coefficient, and $\nu$ is the kinematic viscosity. The external force $\bm{f}$ is assumed to be stationary in time and divergence-free in space. We are interested in a moderately turbulent regime, when the nonlinear term in the Navier-Stokes equation plays a substantial role. Then, despite the stationary nature of the external forcing, the fluid flow is characterized by strong spatiotemporal variations. Now suppose that someone measures the velocity field in a certain region (possibly with open boundaries) and makes observations for several times longer than the typical time of flow fluctuations. In what follows, we will demonstrate how to reconstruct the dense velocity $\bm{v}(\bm{r},t)$ and pressure $p(\bm{r},t)$ fields, and establish the driving force $\bm{f}(\bm{r})$ and unknown fluid viscosity $\nu$ and bottom friction $\alpha$, based only on potentially sparse and noisy velocity field data. Let us emphasize that the developed approach can be easily implemented on top of the existing PIV/PTV workflows, thereby expanding their functionality.
The contribution of this paper compared to previous studies is twofold. First, we generate a novel dataset of forced two-dimensional fluid flow that can serve as a benchmark solution in future research Hao et al. (2023). Second, we extend the PINN algorithm by including the ability to infer driving forces. This is achieved by simultaneously training two neural networks (NNs), one for the velocity and pressure field and the other for the external force. This division, in particular, makes it possible to explicitly take into account the stationarity of the external forcing in the design of the corresponding NN. Including physical information in NN architecture is known to enhance training performance and prediction accuracy.
II Data generation
Data for this study are obtained by integrating (1) using the GeophysicalFlows.jl pseudospectral code Constantinou et al. (2021), which has recently been successfully applied to model two-dimensional turbulence Parfenyev (2022); Kolokolov, Lebedev, and Parfenyev (2024); Parfenyev (2024). The code can be executed on the GPU, resulting in high computational performance. The aliasing errors are removed with the two-third rule. The simulation domain is a doubly periodic box of size $2\pi\times 2\pi$. The boundary conditions and form of the domain are only relevant to the data production process; the rest of the study is generic in this regard.
The observation area is smaller and has the size of $L\times L$ with $L=\pi$, see Fig. 1a. Thus, fluid flows in and out of the observation zone and the movement within it depends on the unseen surroundings. The system parameters $\nu=0.01$ and $\alpha=0.1$ are chosen to be consistent with the experimental studies. The external forcing has the Kolmogorov form
$f_{x}=f_{0}\sin(k_{f}y),\quad f_{y}=0,$ (2)
with $f_{0}=10$ and $k_{f}=5$. We start from an arbitrary random initial condition, and after a transient process, the flow reaches a statistical steady-state. Once stationary, we saved data at $\Delta t=0.02$ intervals for $T=4$ units, covering several turn-over times of flow fluctuations. The spatial resolution is $512^{2}$ for the entire computational domain and $256^{2}$ for the observation zone.
Fig. 1a shows snapshot of the vorticity field $\omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ in the statistical steady-state. The flow is chaotic, with no obvious imprint of an external forcing. Since there are two dissipation mechanisms – viscosity and bottom friction – the flow state can be characterized by two dimensionless parameters Mishra et al. (2015): the Reynolds number $Re=UL/\nu\approx 1.3\times 10^{3}$ and $Rh=U/(\alpha L)\approx 13$, which were computed using the root mean square velocity $U$ and the size $L$ of the observation domain. The energy spectrum is presented in Fig. 1b. It indicates that fluid flow is mainly determined by modes with wave vectors $k\lesssim k_{f}$.
Figure 1: Vorticity field (a) and energy spectrum (b) in a statistical steady-state. The shaded area shows the observation zone.
Next, we assume that measurements reveal the velocity field $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ at $N_{data}$ points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, which are randomly scattered within the observation area and time interval of length $T$. These data imitate PIV/PTV measurements and will be used for subsequent neural network training. To study the influence of measurement noise, we will distort the velocity as $v_{x,y}^{d}\rightarrow(1+\eta)v_{x,y}^{d}$, where $\eta\sim\mathcal{N}(0,\varepsilon^{2})$ is a normally distributed random variable and the standard deviation $\varepsilon$ controls the noise level. We set $N_{data}=3\times 10^{4}$, which is around $0.2\%$ of the total data and thus represents sparse measurements. On average, this corresponds to $150$ points per snapshot compared to $65536$ points at full resolution.
III Physics-Informed Neural Network
In our study, we utilize two independent NNs that are trained simultaneously, see Fig. 2a. The first NN approximates the velocity and pressure at a particular position and time. It has three input neurons and three output neurons, $(v_{x},v_{y},p)=\mathcal{F}_{\vec{\theta}_{1}}(x,y,t)$. The second NN is intended to predict stationary force. As a result, it contains only two input neurons and two output neurons, $(f_{x},f_{y})=\mathcal{G}_{\vec{\theta}_{2}}(x,y)$. The parameters $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$ denote the trainable variables for NNs.
Figure 2: (a) Schematic of the neural networks and (b) adopted residual network architecture.
Figure 3: Comparison between DNS data (ground truth) and PINN predictions for model $A_{1}$ (with clean data).
NNs must be sufficiently complicated to reproduce patterns that are essential for the considered turbulent flow. Universal approximation theorem ensures that this is possible Cybenko (1989); Hornik, Stinchcombe, and White (1989). In the original paper Raissi, Perdikaris, and Karniadakis (2019), the authors used a fully connected NN architecture, but here we implement a residual design He et al. (2016) for both NNs as we found that it performs better in agreement with the results reported earlier Cheng and Zhang (2021). The adopted architecture is shown in Fig. 2b. The parameters $N_{l}$ and $N_{n}$ represent the number of hidden layers and neurons in each hidden layer, respectively. The output $\bm{Y}_{k}$ of $k$-layer is given by
$\bm{Y}_{2j+1}=\sigma\left(\hat{W}_{2j+1}\bm{Y}_{2j}+\bm{b}_{2j+1}+\bm{Y}_{2j-1}\right),\quad j\geq 0,$
$\bm{Y}_{2j}=\sigma\left(\hat{W}_{2j}\bm{Y}_{2j-1}+\bm{b}_{2j}\right),\quad j\geq 1,$ (3)
$\bm{Y}_{N_{l}+1}=\hat{W}_{N_{l}+1}\bm{Y}_{N_{l}}+\bm{b}_{N_{l}+1},$
where $\bm{Y}_{-1}\equiv\bm{0}$ for convenience of notation, $\bm{Y}_{0}$ corresponds to the input layer, $\bm{Y}_{N_{l}+1}$ denotes the output layer, $\hat{W}$ and $\bm{b}$ are weights and biases (trainable parameters that comprise $\vec{\theta}$), and $\sigma$ is the nonlinear activation function applied element-wise. We use a tanh activation function because it gives a reasonable trade-off between training time and resulting accuracy Wang, Liu, and Wang (2022).
In general, increasing the size of NNs can reduce the prediction error, but only when the NNs are given enough data Wang, Liu, and Wang (2022). Here we set $N_{l}=7$ and $N_{n}=250$ for $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$, and $N_{l}=5$ and $N_{n}=30$ for $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$. This choice of parameters is consistent with $N_{data}=3\times 10^{4}$. Note that $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$ is much simpler than $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$ because it is not time dependent and the external forcing usually does not have a complex spatial dependency.
Training NNs involves minimizing a loss function consisting of two terms. The first term represents the deviation between predicted and measured velocity
$\mathcal{L}_{data}=\dfrac{1}{N_{data}U^{2}}\sum_{n=1}^{N_{data}}|\bm{v}(\bm{r}_{n}^{d},t_{n}^{d})-\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})|^{2},$ (4)
where $U^{2}=\sum_{n}\bm{v}_{d}^{2}(\bm{r}_{n}^{d},t_{n}^{d})/N_{data}$ is measurement-based estimate of mean squared velocity. The second term penalizes deviations of predictions from the Navier-Stokes equation and incompressibility conditions
$\mathcal{L}_{eq}=\dfrac{1}{N_{eq}}\sum_{n=1}^{N_{eq}}\left[\bm{e}_{1}^{2}(\bm{r}_{n}^{e},t_{n}^{e})+e_{2}^{2}(\bm{r}_{n}^{e},t_{n}^{e})+e_{3}^{2}(\bm{r}_{n}^{e})\right],$ (5)
where $\{\bm{r}_{n}^{e},t_{n}^{e}\}_{n=1}^{N_{eq}}$ are the collocation points at which equation loss is calculated and
$\bm{e}_{1}=\left(\partial_{t}\bm{v}+(\bm{v}\nabla)\bm{v}+\nabla p+\alpha\bm{v}-\nu\nabla^{2}\bm{v}-\bm{f}\right)L/U^{2},\;$ (6)
$e_{2}=\left(\partial_{x}v_{x}+\partial_{y}v_{y}\right)L/U,$ (7)
$e_{3}=\left(\partial_{x}f_{x}+\partial_{y}f_{y}\right)L^{2}/U^{2}.$ (8)
Computing the right-hand sides involves differentiating the outputs of NNs with respect to input neurons using automatic differentiation Baydin et al. (2018). The number and location of collocation points are not limited by measurements and can be chosen arbitrarily. We set $N_{eq}=3N_{data}$, with one-third of these points corresponding to data points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, and the rest are generated randomly. Note that instead of requiring the incompressibility conditions to be satisfied, the NNs can be modified to predict the relevant stream functions Raissi, Perdikaris, and Karniadakis (2019); Kag, Seshasayanan, and Gopinath (2022). However, in this case, calculating the terms in the Navier-Stokes equation involves computation of higher-order derivatives, which significantly slows down the training process.
Finally, the total loss function is given by
$\mathcal{L}=\mathcal{L}_{data}+\beta\mathcal{L}_{eq},$ (9)
where $\beta$ is a weighting coefficient used to balance two terms. Its value can be adaptively changed during training Wang, Teng, and Perdikaris (2021); Jin et al. (2021), but here we use a simpler strategy in which the weight is fixed. Empirically, we have found that $\beta=0.02$ is well suited for this type of problem.
Minimization of the loss function (9) is performed by a gradient-based optimizer that adjusts network and model parameters. As the bottom friction and fluid viscosity are positive, it is convenient to represent them in the form $\alpha=e^{\lambda_{1}}$ and $\nu=e^{\lambda_{2}}$. So, the training variables are $\{\vec{\theta}_{1},\vec{\theta}_{2},\lambda_{1},\lambda_{2}\}$. First, we train the model for $10^{5}$ epochs using Adam optimizer Kingma and Ba (2014) with learning rate of $\mu=10^{-3}$. To reduce computation time, we have implemented a mixed precision technique Micikevicius et al. (2017). Then we use L-BFGS-B optimizer Byrd et al. (1995) until the tolerance reaches machine precision. The training is carried out on a single NVIDIA A100 GPU. The code is written using the PyTorch library Paszke et al. (2019) and is openly available at [rep, ].
Figure 4: Velocity (left), pressure (middle) and force (right) probability density functions for DNS data and PINN models trained with various noise levels in the input data. All figures share the same legend. The dashed line corresponds to expression (11).
IV Results
[…]
$\Delta_{Q}=\left\langle\dfrac{||Q-Q_{DNS}||_{2}}{||Q_{DNS}||_{2}}\right\rangle_{T},$ (10)
where $Q$ denotes the physical quantity of interest predicted by the PINN, and $Q_{DNS}$ is its exact value. Table 1 shows the results for velocity, pressure and external force, as well as the relative errors in determining the bottom friction $\alpha$ and fluid viscosity $\nu$. The developed method consistently reconstructs the velocity field in the observation region with an accuracy of about $0.2\%$, while the model parameters $\alpha$ and $\nu$ are determined with an error of several percent. Relative errors in predicting the pressure field and external force are larger, with greater dispersion and an obvious correlation between them. The average for both is about $8\%$, and the standard deviation is close to $5\%$. This behavior is due to the lack of measurements for these quantities, so the model extracts them solely from the analysis of physical laws.
Table 1: Relative $L_{2}$ errors in predicting velocity, pressure, external force, bottom friction and fluid viscosity (with clean data).
model $\Delta_{\bm{v}},\%$ $\Delta_{p-\bar{p}},\%$ $\Delta_{\bm{f}},\%$ $\Delta_{\alpha},\%$ $\Delta_{\nu},\%$
$A_{0}$ $0.22$ $5.52$ $6.09$ $3.66$ $1.42$
$A_{1}$ $0.21$ $4.21$ $4.02$ $4.27$ $1.56$
$A_{2}$ $0.20$ $3.11$ $4.10$ $2.83$ $1.41$
$A_{3}$ $0.23$ $17.34$ $16.58$ $4.36$ $1.50$
$A_{4}$ $0.20$ $2.13$ $2.28$ $3.22$ $1.26$
$A_{5}$ $0.24$ $8.34$ $8.39$ $3.84$ $1.76$
$A_{6}$ $0.22$ $10.46$ $10.17$ $4.02$ $1.47$
$A_{7}$ $0.21$ $17.37$ $16.59$ $4.21$ $1.53$
$A_{8}$ $0.21$ $3.85$ $3.66$ $4.07$ $1.65$
$A_{9}$ $0.23$ $7.48$ $7.25$ $4.10$ $1.59$
Table 2: Dependence of mean values and standard deviations for prediction errors on the noise level.
noise $\overline{\Delta}_{\bm{v}},\%$ $\overline{\Delta}_{p-\bar{p}},\%$ $\overline{\Delta}_{\bm{f}},\%$ $\overline{\Delta}_{\alpha},\%$ $\overline{\Delta}_{\nu},\%$
$0\%$ $0.22\pm 0.01$ $8.0\pm 5.6$ $7.9\pm 5.2$ $3.9\pm 0.5$ $1.5\pm 0.1$
$0.5\%$ $0.28\pm 0.01$ $8.9\pm 5.4$ $8.6\pm 5.1$ $3.1\pm 0.9$ $1.2\pm 0.2$
$1\%$ $0.50\pm 0.03$ $7.8\pm 5.9$ $7.6\pm 5.6$ $2.2\pm 1.8$ $0.7\pm 0.5$
$2\%$ $2.39\pm 0.06$ $12.9\pm 5.1$ $12.2\pm 5.3$ $41.4\pm 4.4$ $12.3\pm 0.9$
$5\%$ $5.84\pm 0.28$ $19.1\pm 5.5$ $17.9\pm 6.0$ $188\pm 26$ $56.1\pm 2.0$
Next, we repeat the training, but now the models receive noisy measurement data $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$. Table 2 reports the dependence of the average values and standard deviations for prediction errors on the noise level, obtained by averaging over ten models in each case. The developed algorithm is robust to small noise ($\leq 1\%$) in the input data. Moreover, the error for the reconstructed velocity field in this case is less than the noise level. Thus, PINN is able to correct for small noise based on the knowledge that the velocity field must satisfy the Navier-Stokes equation and incompressibility condition. As the noise level increases ($\geq 2\%$), the performance of the trained models gradually decreases. Since for a turbulent flow the dissipative terms make a small contribution to the Navier-Stokes equation, the model quickly loses the ability to predict the bottom friction $\alpha$ and fluid viscosity $\nu$. At the same time, the accuracy of the velocity field determination remains comparable to the noise level. Errors in predicting the pressure field and external force are larger and amount to tens of percent.
To characterize the statistical properties of reconstructed dense velocity, pressure and force fields, we calculate the corresponding probability density functions (PDFs); see Fig. 4. To illustrate, we used trained models with pressure and force errors close to the mean values presented in Table 2. The velocity PDFs are almost identical to the DNS data even for PINN models based on noisy input data, confirming the low $L_{2}$ error between them. The pressure field is reconstructed worse, and the main difference from the DNS data lies in the typical values, and not in the tails of the distribution function. The force PDF corresponding to expression (2) can be calculated analytically
$\rho(f)=\dfrac{2}{\pi f_{0}}\left(1-\dfrac{f^{2}}{f_{0}^{2}}\right)^{-1/2},$ (11)
where $f=(f_{x}^{2}+f_{y}^{2})^{1/2}\leq f_{0}$, and it diverges at $f\to f_{0}$. The largest disparities between the reconstructed PDFs and the DNS data are concentrated around this point.
Figure 5: Energy spectra for DNS data and PINN models trained with various noise levels in the input data.
In Fig. 5, we compare the energy spectra of the reference flow and PINN reconstructions, which are commonly used to characterize the statistical properties of a velocity field. Since the observation region is not periodic, when constructing the spectra we applied a Hann window function. For this reason, the DNS spectrum is slightly different from that presented earlier in Fig. 2b. As expected from the previous analysis, the overall errors are small. They are mainly concentrated on small scales with $k\gg k_{f}$ and gradually rise with increasing noise level in the input data.
V Conclusion
We showed that PINN can combine measurement data with an accurate physical model to extract additional information from ordinary flow observations. The developed technique can be built on top of current PIV/PTV workflows, potentially improving their functionality. In particular, we investigated a forced two-dimensional turbulent fluid flow and demonstrated that the proposed method can reconstruct the dense velocity and pressure fields, infer the driving force, and determine the unknown fluid viscosity and bottom friction coefficient using only sparse and noisy velocity measurements.
The suggested technique trains two independent NNs simultaneously, one to estimate the velocity and pressure field and the other to approximate the external forcing. This enables us to use different designs in NNs, which makes it possible to take into account the physical features of the considered system. In our case, this allowed us to take into account the stationarity of the external forcing implemented in the physical model. Training of NNs involves minimization of the loss function that penalizes deviations from measured data and violations of the Navier-Stokes equation and incompressibility conditions. While our intention was to select the optimal NN design and hyperparameter configuration, a thorough examination of this issue was outside the scope of this work. We anticipate that the introduction of more advanced NN architectures Zhao, Ding, and Prakash (2023); Wang et al. (2024); Liu et al. (2024); Buzaev et al. (2023) can further improve the developed method.
The reconstructed flows were compared to the reference flow in terms of $L_{2}$ errors, PDFs of the reconstructed fields, and energy spectra. Overall, the results show good accuracy for super-resolution and inference applications and demonstrate moderate robustness to noise in the input data. At low noise levels, the algorithm can correct errors in the velocity field measurements based on information from the physical model. The developed technique can be useful for analyzing flow observations when the physical model of the phenomenon is well constructed.
Acknowledgements.
The work of VP and IN was supported by the Ministry of Science and Higher Education of the Russian Federation within the State Assignment No. FFWR-2024-0017 of Landau Institute for Theoretical Physics. VP also acknowledges support from the Foundation for the Advancement of Theoretical Physics and Mathematics ”BASIS”, Project No. 22-1-3-24-1. The authors are grateful to the Landau Institute for Theoretical Physics for providing computing resources.
Data AVAILABILITY
The data that support the findings of this study are openly available at https://github.com/parfenyev/2d-turb-PINN/.